# Ratio asymptotics and zero density for orthogonal polynomials with varying Verblunsky coefficients

### v1

@@ 3. Ratio asymptotics for varying POPUC and OPUC
$\Phi_{n,N}(z)=\det\left(z-\mathcal{C}_{n}(\alpha_{0,N},\alpha_{1,N},\dots,\alpha_{n-1,N})\right)$ (3.1)
@@
$\Phi_{n,N}^{(\beta)}(z)=\det\left(z-\mathcal{C}_{n}(\alpha_{0,N},\alpha_{1,N},\dots,\alpha_{n-2,N},\beta)\right),$ (3.2)
@@ 3.1. Asymptotically constant Verblunsky coefficients
For the formulation of the ratio asymptotic formulas, we need to introduce the sets of limit points of zeros of POPUC. Let $\Lambda_{\beta}$ denote the set of (weak) limit points of zeros of POPUC $\Phi_{n,N}^{(\beta)}$, i.e.,
@@
$\Sigma_{n,N}^{(\beta)}=\big\{z\in{\partial\mathbb{D}}\;\big|\;\Phi_{n,N}^{(\beta)}(z)=0\big\}.$
@@
$\lim_{n/N\to t}|\alpha_{n,N}|=a,$ (3.12)
$\lim_{n/N\to t}\alpha_{n+1,N}/\alpha_{n,N}=\lambda,$ (3.13)
@@ 3.2. Asymptotically periodic Verblunsky coefficients
$\mathcal{Z}(f)$ $=\{z\in{\mathbb{D}}\mid f(z)=0\},$ (3.15)
$\mathcal{P}(f)$ $=\{z\in{\mathbb{C}}\setminus\overline{{\mathbb{D}}}\mid f(z)=\infty\},$ (3.16)
for the sets of zeros and poles of a Schur function $f$. By analyticity of $f$ on $\mathbb{D}$} and (2.13), $\mathcal{Z}(f)$ and $\mathcal{P}(f)$ are indeed subsets of $\mathbb{D}$} and ${\mathbb{C}}\setminus\overline{{\mathbb{D}}}$, respectively.
@@
It is clear that the same argument can be extended to the setting of right limits, using essentially the same proof (in particular, one can allow $\alpha_{k}\in\overline{{\mathbb{D}}}$). We do not pursue this here. See [BRZ] for a related setting.
@@
The only difference in the proof of (ii) is that, unlike Theorem 5(ii), where $m_{-\bar{\alpha}\beta}(\bar{z}^{-1})$ never vanishes on ${\mathbb{C}}\setminus\Lambda_{\beta,t}$, in our case $m$ may vanish. We therefore need to avoid points $z$ where $m_{\beta}^{(j)}(\bar{z}^{-1})=0$. By Lemma 1(ii) and (2.13), these points are exactly those in $\mathcal{P}(f_{\beta}^{(j)})$. This set is $\beta$-independent, so we may take $\beta=1$, and the proof of (ii) is complete.
@@
$\frac{\Phi_{n+1,N}^{(\beta)}(z)}{\Phi_{n,N}^{(\beta)}(z)}=\frac{\Phi_{n+1,N}^{(\beta)}(z)}{\Phi^{*}_{n,N}(z)}\frac{\Phi^{*}_{n-1,N}(z)}{\Phi_{n,N}^{(\beta)}(z)}\frac{\Phi^{*}_{n,N}(z)}{\Phi^{*}_{n-1,N}(z)},$ (3.19)
and arguing as in (3.9) while applying (i) and (iii), uniformly on compacts of ${\mathbb{C}}\setminus(\Lambda_{\beta,t}\cup\mathcal{P}(f^{(j-1)}_{1}))$. Since $\mathcal{Z}(f^{(j-1)}_{1}))$ belongs to $\mathbb{D}$} and $\mathcal{P}(f^{(j-1)}_{1})$ to ${\mathbb{C}}\setminus\overline{{\mathbb{D}}}$, we obtain uniform convergence on compacts of ${\mathbb{C}}\setminus\Lambda_{\beta,t}$, as claimed. ∎
Barrios and Lopez [BarLop, Thm. 2], using a very different proof, computed the limit of the ratios $\Phi_{n+p}(z)/\Phi_{n}(z)$. This result can be recovered easily as a corollary of the previous theorem.
@@
$\mathcal{Z}$ $=\bigcup_{j=0}^{p-1}\mathcal{Z}(f^{(j)}_{1})$, (3.21)
$\mathcal{P}$ $=\bigcup_{j=0}^{p-1}\mathcal{P}(f^{(j)}_{1})$. (3.22)
@@
Notice that $L(z)$ is fully determined by the sequence $(\alpha_{j})_{j=0}^{p-1}$. So without loss of generality we can assume that $\alpha_{kp+j,N}=\alpha_{j}$ for all $k,N\in\mathbb{N}_{0}$ and $j=0,\ldots,p-1$. It is an elementary calculus exercise to show that if a sequence $\{x_{k}\}$ satisfies $\lim x_{k+1}/x_{k}=a$ and $\lim|x_{k}|^{1/k}=b$ then $|a|=b$. Applying this to $x_{k}=\Phi_{kp}(z)$ and using (2.47) with (3.23), we obtain
@@
Now note that $L(z)$ in (3.23) is analytic nonvanishing function on $\mathbb{D}$} (recall that $|f_{1}^{(j-1)}(z)|\leq 1$ for Schur functions). $G_{\Delta}$ is analytic on $\mathbb{D}$} and it is also nonvanishing, by (3.24). Finally, $L(0)=G_{\Delta}(0)=1$ by (3.23) and (2.41). Then applying the maximum modulus principle to $L(z)/G_{\Delta}(z)$ shows that $L(z)=G_{\Delta}(z)$ and completes the proof. ∎
@@ 4. Asymptotic zero density for variable POPUC
in (4.2) is an average of the equilibrium measures $\nu_{|\alpha(s)|}$ over $s\in[0,t]$, supported on $\{e^{{\rm i}\theta}\mid\theta_{|\alpha^{*}(t)|}\leq\theta\leq 2\pi-\theta_{|\alpha^{*}(t)|}\}$, where
@@
$\frac{{\rm d}\sigma_{t}}{{\rm d}\theta}\left(e^{{\rm i}\theta}\right)=\frac{1}{2\pi t}\int_{0}^{t}\frac{\sin\left(\theta/2\right)}{\sqrt{\sin^{2}\left(\theta/2\right)-|\alpha(s)|^{2}}}\,\chi_{\left[\theta_{|\alpha(s)|},2\pi-\theta_{|\alpha(s)|}\right]}(\theta){\rm d}s,$ (4.4)
@@ 4.1. First proof based on ratio asymptotics and potential-theoretic argument
By Widom’s lemma, if $U_{\mu_{n}}(z)$ tends to a limiting function $U(z)$, as $n\to\infty$, for almost every $z\in{\mathbb{C}}$ (with respect to the two-dimensional Lebesgue measure in $\mathbb{C}$}), then $\mu_{n}$ converges weakly to a measure $\mu$, as $n\to\infty$, and $U$ is the logarithmic potential of $\mu$. The limiting measure $\mu$ can be determined from $U$ by the formula $2\pi\mu=\Delta U$ which is to be understood in the distributional sense.
@@
$\frac{\Phi_{n}(z)}{\Phi_{n}^{(\beta)}(z)}=\frac{z-\bar{\alpha}_{n-1}\Phi_{n-1}^{*}(z)\big/\Phi_{n-1}(z)}{z-\bar{\beta}\Phi_{n-1}^{*}(z)\big/\Phi_{n-1}(z)},$ (4.10)
@@
Note that the above sum is the Riemann sum for a piece-wise constant function, i.e., we have
@@ 4.2. Second proof based on method of moments
We assume that the collection of matrices $\{a_{j,k}^{(N)}\mid j,k\in{\mathbb{Z}},|j-k|\leq 1,N\in\mathbb{N}_{0}\}$ is uniformly bounded, so that $J^{(N)}$, regarded as an operator on $\ell^{2}({\mathbb{Z}})$, is bounded.
Let $P_{n}$ be the orthogonal projection in $\ell^{2}({\mathbb{Z}})$ onto the span of $\{e_{j}\mid 0\leq j\leq n-1\}$. Denote $J_{n}^{(N)}$ to be the truncation $J_{n}^{(N)}=P_{n}J^{(N)}P_{n}{\upharpoonright\mathop{\mathrm{Ran}}\nolimits P_{n}}$. Denote further by $\nu_{n,N}$ the eigenvalue counting measure of $J_{n}^{(N)}$, i.e.
$\nu_{n,N}=\frac{1}{n}\sum_{j=1}^{n}\delta_{z_{j}},$ (4.15)
where $z_{j}$, $k=1,\dots,n$, are the eigenvalues of $J_{n}^{(N)}$ counted with multiplicity.
For three fixed complex $\ell\times\ell$ matrices $\gamma_{-1},\gamma_{0},\gamma_{1}$, let $J(\gamma_{-1},\gamma_{0},\gamma_{1})$ be the tridiagonal block (Laurent) matrix of the form (4.14) with $a_{j-1,j}^{(N)}=\gamma_{-1}$, $a_{j,j}^{(N)}=\gamma_{0}$, $a_{j+1,j}^{(N)}=\gamma_{1}$ for all $j\in{\mathbb{Z}}$. Let $\nu_{n}^{(\gamma_{-1},\gamma_{0},\gamma_{1})}$ be the eigenvalue counting measure of the $n\times n$ truncation of $J(\gamma_{-1},\gamma_{0},\gamma_{1})$ as above. The existence of the weak limit of $\nu_{n}^{(\gamma_{-1},\gamma_{0},\gamma_{1})}$ as $n\to\infty$ need not be guaranteed in general. In the next statement, we assume a weaker condition that the nonnegative integer moments of $\nu_{n}^{(\gamma_{-1},\gamma_{0},\gamma_{1})}$ converge to the corresponding moments of some probability measure on $\mathbb{C}$}.
@@
For $k=0$ the claim is trivial. Suppose $k\in\mathbb{N}$. First note that the left-hand side of (4.18) is equal to
$\lim_{n/N\to t}\frac{1}{n}\mathop{\mathrm{Tr}}\nolimits\left(P_{n}J^{(N)}P_{n}\right)^{k}=\lim_{n/N\to t}\frac{1}{n}\mathop{\mathrm{Tr}}\nolimits P_{n}\left(J^{(N)}\right)^{k}P_{n}.$ (4.19)
We denote
@@
$\int_{0}^{1}\mathop{\mathrm{Tr}}\nolimits\sum_{(d_{1},\dots,d_{k})\in\mathcal{D}_{k}}a_{[ns],[ns]+d_{1}}^{(N)}a_{[ns]+d_{1},[ns]+d_{1}+d_{2}}^{(N)}\dots a_{[ns]+d_{1}+\dots+d_{k-1},[ns]}^{(N)}\,{\rm d}s.$ (4.20)
@@
where $\mathcal{C}_{n,N}^{(\beta)}=\mathcal{C}_{n}(\alpha_{0,N},\ldots,\alpha_{n-2,N},\beta)$ is our finite $n\times n$ unitary CMV matrix (see (2.5)) acting on $\mathop{\mathrm{Ran}}\nolimits P_{n}$; the first zero matrix acts on $\ell^{2}(-\mathbb{N})$, and the second on $\ell^{2}(\{n,n+1,\ldots\})$. By the assumption (4.1) and (2.6), we obtain the existence (for a.e. $s\in[0,t]$) of limits (4.16) with
$\gamma_{-1}(s)={\small\begin{pmatrix}0&\overline{\alpha(s)}\rho(s)\\ 0&\rho(s)^{2}\end{pmatrix}},\quad\gamma_{0}(s)={\small\begin{pmatrix}-|\alpha(s)|^{2}&\overline{\alpha(s)}\rho(s)\\ -{\alpha(s)}\rho(s)&-|\alpha(s)|^{2}\end{pmatrix}},\quad\gamma_{1}(s)={\small\begin{pmatrix}\rho(s)^{2}&0\\ -{\alpha(s)}\rho(s)&0\end{pmatrix}}$,
where $\rho(s)=(1-|\alpha(s)|^{2})^{1/2}$.

### v2

@@ 3. Ratio asymptotics for varying POPUC and OPUC
$\Phi_{n,N}(z)=\det\left(z-\mathcal{C}_{n}(\alpha_{0,N},\alpha_{1,N},\dots,\alpha_{n-1,N})\right)$
@@
$\Phi_{n,N}^{(\beta)}(z)=\det\left(z-\mathcal{C}_{n}(\alpha_{0,N},\alpha_{1,N},\dots,\alpha_{n-2,N},\beta)\right),$
@@ 3.1. Asymptotically constant Verblunsky coefficients
For the formulation of the ratio asymptotic formulas, we need to introduce the sets of limit points of zeros of POPUC. Let $\Lambda_{\beta}$ denote the set of (weak) limit points of zeros of $\Phi_{n,N}^{(\beta)}$, i.e.
@@
$\mathcal{Z}(\Phi_{n,N}^{(\beta)})=\big\{z\in{\partial\mathbb{D}}\;\big|\;\Phi_{n,N}^{(\beta)}(z)=0\big\}.$
@@
$\lim_{n/N\to t}|\alpha_{n,N}|=a\quad\mbox{ and }\quad\lim_{n/N\to t}\frac{\alpha_{n+1,N}}{\alpha_{n,N}}=\lambda,$
@@ 3.2. Asymptotically periodic Verblunsky coefficients
Note that here we assume that $\alpha_{j}\in{\mathbb{D}}$, that is, we exclude the case when some of $\alpha_{j}$’s are unimodular. This is done to simplify the exposition. With enough care this case can be easily accommodated using the same methods.
@@
$\mathcal{Z}(f)$ $=\{z\in{\mathbb{D}}\mid f(z)=0\},$
$\mathcal{P}(f)$ $=\{z\in{\mathbb{C}}\setminus\overline{{\mathbb{D}}}\mid f(z)=\infty\},$
for the sets of zeros and poles of a Schur function $f$. By analyticity of $f$ on $\mathbb{D}$} and (2.12), the set of zeros and poles of $f$ are indeed subsets of $\mathbb{D}$} and ${\mathbb{C}}\setminus\overline{{\mathbb{D}}}$, respectively.
@@
We remark that the same argument applies to the setting of right limits, using essentially the same proof (in particular, one can allow $\alpha_{k}\in\overline{{\mathbb{D}}}$). We do not pursue this here. See [BRZ] for a related setting.
@@
The only difference in the proof of (ii) is that, unlike in Theorem 5(ii), where $m_{-\bar{\alpha}\beta}(\bar{z}^{-1})$ never vanishes on ${\mathbb{C}}\setminus\Lambda_{\beta,t}$, in our case the $m$-function may vanish. We therefore need to avoid points $z$ where $m_{\beta}^{(j)}(\bar{z}^{-1})=0$. By Lemma 1(ii) and (2.12), these points are exactly those in $\mathcal{P}(f_{\beta}^{(j)})$. This set is $\beta$-independent, so we may take $\beta=1$, and the proof of (ii) is complete.
@@
$\frac{\Phi_{n+1,N}^{(\beta)}(z)}{\Phi_{n,N}^{(\beta)}(z)}=\frac{\Phi_{n+1,N}^{(\beta)}(z)}{\Phi^{*}_{n,N}(z)}\frac{\Phi^{*}_{n-1,N}(z)}{\Phi_{n,N}^{(\beta)}(z)}\frac{\Phi^{*}_{n,N}(z)}{\Phi^{*}_{n-1,N}(z)},$
and arguing as in (3.7) while applying (i) and (iii), uniformly on compacts of ${\mathbb{C}}\setminus(\Lambda_{\beta,t}\cup\mathcal{P}(f^{(j-1)}_{1}))$. Since $\mathcal{Z}(f^{(j-1)}_{1})$ belongs to $\mathbb{D}$} and $\mathcal{P}(f^{(j-1)}_{1})$ to ${\mathbb{C}}\setminus\overline{{\mathbb{D}}}$, we obtain uniform convergence on compacts of ${\mathbb{C}}\setminus\Lambda_{\beta,t}$, as claimed. ∎
Barrios and López [BarLop, Thm. 2], using a very different proof, computed the limit of the ratios $\Phi_{n+p}(z)/\Phi_{n}(z)$. This result can be recovered easily as a corollary of the previous theorem.
@@
$\mathcal{Z}=\bigcup_{j=0}^{p-1}\mathcal{Z}(f^{(j)}_{1})\quad\mbox{ and }\quad\mathcal{P}=\bigcup_{j=0}^{p-1}\mathcal{P}(f^{(j)}_{1})$.
@@
Notice that $L(z)$ is fully determined by the sequence $(\alpha_{j})_{j=0}^{p-1}$. So without loss of generality we can assume that $\alpha_{kp+j,N}=\alpha_{j}$ for all $k,N\in\mathbb{N}_{0}$ and $j=0,\ldots,p-1$. It is an elementary calculus exercise to show that if a complex sequence $\{x_{k}\}$ satisfies $x_{k+1}/x_{k}\to a$ and $|x_{k}|^{1/k}\to b$, as $k\to\infty$, then $|a|=b$. Applying this to $x_{k}=\Phi_{kp}(z)$ and using (2.35) with (3.14), we obtain
@@
Now note that $L(z)$ in (3.14) is analytic nonvanishing function on $\mathbb{D}$}, bearing in mind that $|f_{1}^{(j-1)}(z)|\leq 1$ for Schur functions. Function $G_{\Delta}$ is analytic on $\mathbb{D}$} and it is also nonvanishing, by (3.15). Finally, $L(0)=G_{\Delta}(0)=1$ by (3.14) and (2.33). Then applying the maximum modulus principle to $L/G_{\Delta}$ shows that $L=G_{\Delta}$ and completes the proof. ∎
@@ 4. Asymptotic zero density for variable POPUC
in (4.2) is an average of the equilibrium measures $\nu_{|\alpha(s)|}$ over $s\in[0,t]$ supported on the arc $\{{\rm e}^{{\rm i}\theta}\mid\theta_{|\alpha^{*}(t)|}\leq\theta\leq 2\pi-\theta_{|\alpha^{*}(t)|}\}$, where
@@
$\frac{{\rm d}\sigma_{t}}{{\rm d}\theta}\left({\rm e}^{{\rm i}\theta}\right)=\frac{1}{2\pi t}\int_{0}^{t}\frac{\sin\left(\theta/2\right)}{\sqrt{\sin^{2}\left(\theta/2\right)-|\alpha(s)|^{2}}}\,\chi_{\left[\theta_{|\alpha(s)|},2\pi-\theta_{|\alpha(s)|}\right]}(\theta){\rm d}s,$ (4.4)
@@ 4.1. First proof based on ratio asymptotics and potential-theoretic argument
By Widom’s lemma, if $U_{\mu_{n}}(z)$ tends to a limiting function $U(z)$, as $n\to\infty$, for a.e. $z\in{\mathbb{C}}$ (with respect to the two-dimensional Lebesgue measure in $\mathbb{C}$}), then $\mu_{n}$ converges weakly to a measure $\mu$, as $n\to\infty$, and $U$ is the logarithmic potential of $\mu$. The limiting measure $\mu$ can be determined from $U$ by the formula $2\pi\mu=\Delta U$ which is to be understood in the distributional sense.
@@
$\frac{\Phi_{n}(z)}{\Phi_{n}^{(\beta)}(z)}=\frac{z-\bar{\alpha}_{n-1}\Phi_{n-1}^{*}(z)\big/\Phi_{n-1}(z)}{z-\bar{\beta}\Phi_{n-1}^{*}(z)\big/\Phi_{n-1}(z)},$
@@
Note that the above sum is the Riemann sum for a piece-wise constant function, i.e.
@@ 4.2. Second proof based on method of moments
We assume that the collection of matrices $\{a_{j,k}^{(N)}\mid j,k\in{\mathbb{Z}},|j-k|\leq 1,N\in\mathbb{N}_{0}\}$ is uniformly bounded. Then $J^{(N)}$, regarded as an operator on $\ell^{2}({\mathbb{Z}})$, is bounded.
Let $P_{n}$ be the orthogonal projection in $\ell^{2}({\mathbb{Z}})$ onto the span of $\{e_{j}\mid 0\leq j\leq n-1\}$. Denote $J_{n}^{(N)}=P_{n}J^{(N)}P_{n}{\upharpoonright\mathop{\mathrm{Ran}}\nolimits P_{n}}$, i.e., $J_{n}^{(}N)$ is the $n\times n$ submatrix of $J^{(N)}$ corresponding to the row and column indices $j=0,\dots,n-1$. Denote further by $\nu_{n,N}$ the eigenvalue counting measure of $J_{n}^{(N)}$, i.e.
$\nu_{n,N}=\frac{1}{n}\sum_{j=1}^{n}\delta_{z_{j}},$
where $z_{1},\dots,z_{n}$ are the eigenvalues of $J_{n}^{(N)}$ counted with multiplicity.
For three fixed complex $\ell\times\ell$ matrices $\gamma_{-1},\gamma_{0},\gamma_{1}$, let $J(\gamma_{-1},\gamma_{0},\gamma_{1})$ be the tridiagonal block (Laurent) matrix of the form (4.13) with $a_{j-1,j}^{(N)}=\gamma_{-1}$, $a_{j,j}^{(N)}=\gamma_{0}$, $a_{j+1,j}^{(N)}=\gamma_{1}$ for all $j\in{\mathbb{Z}}$. Let $\nu_{n}^{(\gamma_{-1},\gamma_{0},\gamma_{1})}$ be the eigenvalue counting measure of the $n\times n$ truncation $J_{n}^{(N)}$ of $J(\gamma_{-1},\gamma_{0},\gamma_{1})$ as above. The existence of the weak limit of $\nu_{n}^{(\gamma_{-1},\gamma_{0},\gamma_{1})}$ as $n\to\infty$ need not be guaranteed in general. In the next statement, we assume a weaker condition that the nonnegative integer moments of $\nu_{n}^{(\gamma_{-1},\gamma_{0},\gamma_{1})}$ converge to the corresponding moments of a probability measure on $\mathbb{C}$}.
@@
For $k=0$ the claim is trivial. Suppose $k\in\mathbb{N}$. First note that
$\operatorname*{w-lim\ }_{n/N\to t}\int z^{k}{\rm d}\nu_{n,N}=\lim_{n/N\to t}\frac{1}{n}\mathop{\mathrm{Tr}}\nolimits\left(P_{n}J^{(N)}P_{n}\right)^{k}=\lim_{n/N\to t}\frac{1}{n}\mathop{\mathrm{Tr}}\nolimits P_{n}\left(J^{(N)}\right)^{k}P_{n},$ (4.17)
where the second equality is due the banded structure of $J^{(N)}$. Next we denote
@@
$\int_{0}^{1}\mathop{\mathrm{Tr}}\nolimits\sum_{(d_{1},\dots,d_{k})\in\mathcal{D}_{k}}a_{[ns],[ns]+d_{1}}^{(N)}a_{[ns]+d_{1},[ns]+d_{1}+d_{2}}^{(N)}\dots a_{[ns]+d_{1}+\dots+d_{k-1},[ns]}^{(N)}\,{\rm d}s.$
@@
where $\mathcal{C}_{n,N}^{(\beta)}=\mathcal{C}_{n}(\alpha_{0,N},\ldots,\alpha_{n-2,N},\beta)$ is our finite $n\times n$ unitary CMV matrix (see (2.7)–(2.8)) acting on $\mathop{\mathrm{Ran}}\nolimits P_{n}$; the first zero matrix acts on $\ell^{2}(-\mathbb{N})$ and the second on $\ell^{2}(\{n,n+1,\ldots\})$. By the assumption (4.1) and (2.5), we obtain the existence, for a.e. $s\in[0,t]$, of limits (4.14) with
$\gamma_{-1}(s)\!=\!\begin{pmatrix}0&\overline{\alpha(s)}\rho(s)\\ 0&\rho(s)^{2}\end{pmatrix}\!,\;\gamma_{0}(s)\!=\!\begin{pmatrix}-|\alpha(s)|^{2}&\overline{\alpha(s)}\rho(s)\\ -{\alpha(s)}\rho(s)&-|\alpha(s)|^{2}\end{pmatrix}\!,\;\gamma_{1}(s)\!=\!\begin{pmatrix}\rho(s)^{2}&0\\ -{\alpha(s)}\rho(s)&0\end{pmatrix}\!$,
where $\rho(s)=\sqrt{1-|\alpha(s)|^{2}}$.